# Generative AI as a lab partner: a case study

### v1

@@ IV Discussion
The findings in the current study are based on seven lab groups and the work and reflection of, in total, 19 students. As such, they should not be interpreted as directly generalizable. However, this study provides important insights into the potential use, current and future, of GenAI tools in high school science laboratories. First, the findings reveal that the participating students found it practical to have access to a tool like ChatGPT during labs to get answers to simple conceptual questions. Second, groups with less prior knowledge of the underlying physics relevant to fully understanding the lab also struggled with making adequate use of ChatGPT to further their understanding, possibly connected to their inability to gauge the correctness of answers provided by ChatGPT. Here, we discuss the findings in relation to the research questions, as well as the broader implications of using GenAI tools in the high school physics laboratory.
@@
Regarding RQ1: What do students perceive as useful interactions with GenAI tools like ChatGPT during a lab session, our findings agree with previous studies. In agreement with [6], our findings show similarities regarding that students require knowledge about the underlying physics in order for them to gauge whether answers provided by ChatGPT were trustworthy. This is exemplified well by Group 6, one of two groups of third-year students taking the optional course Physics 3, where they could get meaningful assistance from ChatGPT during the analysis stage of the lab in identifying a formula to calculate the speed of sound that included temperature dependence. If not for the group’s prior knowledge of the physics of sound, they might not have been able to conclude that the initial equation provided by ChatGPT was faulty in that the form of the equation was correct, but the gas constant $R$ was given with the wrong units and would have had to be converted prior to using it to calculate the speed of sound. The importance of prior knowledge about the specific area to successfully identify the validity of answers given by ChatGPT is further strengthened by examining an interaction during the lab by Group 4. Although they were one of the groups with second-year students taking Physics 1 and thus having access to very little prior knowledge of the physics of sound, they were identified as performing well during the lab since they were able to identify what data they should collect and perform the analysis without issues. However, they ended up struggling towards the end of the lab session when they were tasked to discuss how acoustic levitation works. Due to their limited to non-existent knowledge of wave mechanics and the physics of sound, they had few tools available to them to challenge what answers ChatGPT gave them. As a result, despite Group 4 demonstrating an understanding of the need to pose not just one question but many to try to identify if the answers provided by ChatGPT made sense or not, they were ultimately led down a path toward a wrong explanation that it is mainly the lower speaker that supplies a force by pushing on the styrofoam particle and thus counteracting gravity. Finally, in most other interactions with ChatGPT during the lab session, the groups seemed to take the answers provided by ChatGPT as simply being the truth. This may be connected to how the students, at times, anthropomorphized ChatGPT during the interviews. Group 3 stated that they might not fully trust the answers if ChatGPT did not provide any sources or citations, and Group 5 stated that ChatGPT can be trusted since it has knowledge about previous research results. In general, during the group interviews before and after the lab, when asked about the capabilities of ChatGPT, students were quick to use words such as knowledge and talk less about it as some tool and more as a person with much knowledge. During the interviews, it was mainly groups 6 and 7 that stood by talking about what ChatGPT can answer and not reliably as dictated by what data it had in its training, thus indicating them being more cognizant of that GenAI tools are not having knowledge and reasoning skills similar to humans.
@@
Concerning RQ2 and RQ3, they focus more on how students interacted with ChatGPT during the lab to overcome specific hurdles. Previous studies provide little information about its capabilities to function as a lab partner. What it does show, however, is that depending on the type of questions and problems students face, it might be able to provide reliable answers, as it can be good at solving subject-specific problems [8, 9]. Further, it has also been identified that GenAI tools, like ChatGPT, can reproduce language and provide answers relevant to tackling problems in laboratory settings [10]. Our findings indicate similar conclusions. In RQ2, How do high school students actively engage with ChatGPT to solve problems during a lab investigating the speed of sound using acoustic levitation, Fig. 3 show that students tend to make use of ChatGPT after having done all or most of the data collection. Together with data from the group interviews where students stressed the major use case of ChatGPT being to answer conceptual questions or to help clarify wordings in the lab instructions, this indicates that despite having little to no knowledge about the phenomenon under study, the groups had little problem collecting data using the experimental setup.
@@
On the other hand, the part the groups did make use of ChatGPT during the lab session was primarily during the analysis and to get help with the discussion questions about the phenomenon of acoustic levitation. Here, the primary beneficial use case for the groups was related to getting help with identifying what equations to make use of, as well as getting an introduction to acoustic levitation. A noteworthy special case was Group 5, who “outsourced” the analysis to ChatGPT after failing to understand what to do with the data by simply asking what the sound speed was at the room’s current temperature. During the analysis, it is important to stress that the lab instructions were intentionally vague in that no equations were provided. Instead, the assumption was that students would utilize a combination of ChatGPT and their formula book to identify how to make use of the data to get the speed of sound using the equation $v=f\lambda$ and compare that to the speed of sound they get when using an equation that included temperature dependence. Additionally, if the groups had time, they were tasked to discuss questions related to acoustic levitation, including the underlying principles of acoustic levitation. As such, it was expected that all groups, regardless of their natural intention to use or not use ChatGPT during the study, were incentivized to at least try to use it during the analysis and discussion part.
@@
Overall, the findings presented in the current study demonstrate fruitful and problematic interactions with ChatGPT in terms of aiding students in solving problems or helping them understand the phenomenon of acoustic levitation. Groups 4 and 6 were both able to use ChatGPT to aid them during the analysis in that Group 4 successfully utilized it to aid them in reasoning about whether the wavelength they got from the data collection was reasonable. Further, both groups showed proficiency in discussing how to identify if the equations given to them from ChatGPT were reliable by drawing upon their prior knowledge of dimensional analysis. Contrary to these, groups 1, 2, 3, and 5 all demonstrated that lacking knowledge about the relevant physics and how GenAI tools like ChatGPT “reasons” or has “knowledge”, having access to such tools in a classroom setting, including laboratories, might end up disadvantageous for the students learning. The reasoning here is that if the teacher is not aware of the student’s questions and problems, they may rely on answers from a GenAI tool that are not disciplinarily accepted and could strengthen problematic interpretations of various concepts. Thus, although ChatGPT can be a helpful tool in the physics laboratory, it is crucial to recognize teachers’ role in identifying each student’s needs in relation to their specific problems and questions. It can be reasonably argued that GenAI tools can serve as valuable discussion partners in the physics lab, provided students are aware of their capabilities and limitations and can effectively prompt them. This has the potential to distribute the teachers’ workload more equitably among students, with those who require additional support able to benefit from the tool.
V Conclusion and outlook
@@
In the current study, we address the research gap related to the need to explore how GenAI tools, specifically the free version ChatGPT at the time of the data collection, could be used in the educational physics laboratory. Our findings are in agreement with previous research regarding the importance of educating students about the capabilities and limitations of AI tools, as well as gaining some level of understanding about how they are developed and “think”. In STEM education specifically, there is a need to further explore and develop educational strategies and material that can help students and teachers better understand how to best utilize and think about the role of GenAI in education, not the least in laboratory settings. As these tools continue to be developed, their capabilities to simulate a trustworthy chain of reasoning will improve, furthering the need to carefully consider what implications interactions with such tools may have when it comes to learning concepts in disciplinary-accepted ways. Here, the role and knowledge of the teacher about AI is of utmost importance. Finally, we have identified that there is a potential to make use of GenAI tools like ChatGPT in the physics laboratory as a lab partner, but that the usefulness looks to be dependent on students knowledge about the phenomenon under study as well as their knowledge about how GenAI tools function. Furthermore, the specific ways in which such a tool is best employed in the physics laboratory require more research and will also be dictated by the future development of such tools.
@@ Acknowledgements.
Sebastian: Conceptualization, Data curation (lead), Formal analysis, Investigation (lead), Methodology, Project administration, Resources (lab design; supporting), Validation (lead), Visualization (lead), Writing - original draft, Writing - review & editing (lead). Andreas: Investigation (supporting), Resources (lab design; lead), Validation (supporting), Visualization (supporting), Writing - original draft (supporting), Writing - review & editing (supporting). Jonas: Data curation (supporting), Investigation (supporting), Supervision, Validation (supporting), Writing - review & editing (supporting).

### v2

@@ VI Discussion
The findings in the current study are based on seven lab groups and the work and reflection of, in total, 19 students. As such, they should not be interpreted as directly generalizable. However, this study provides important insights into the potential use, current and future, of GenAI tools in high school science laboratories. Here, we discuss the findings in relation to the research questions, as well as broader implications of using GenAI tools in laboratory work. The discussion is distilled into four parts: The role of prior knowledge and trustworthiness, Utilizing ChatGPT as a lab partner or a source of information, GenAI as a tool to develop problem-solving skills in laboratory settings, and The ongoing problem with GenAI and hallucination.
VI.1 The role of prior knowledge and trustworthiness
Regarding RQ1: What do students perceive as useful interactions with GenAI tools like ChatGPT during a lab session, our findings find that, in agreement with previous research [6], students require knowledge about the underlying physics in order for them to gauge whether answers provided by ChatGPT were trustworthy. This is exemplified well by Group 6, one of two groups of third-year students taking the optional course Physics 3, where they could get meaningful assistance from ChatGPT during the analysis stage of the lab in identifying a formula to calculate the speed of sound that included temperature dependence. If not for the group’s prior knowledge of the physics of sound, they might not have been able to conclude that the initial equation provided by ChatGPT was faulty in that the form of the equation was correct, but the gas constant $R$ was given with the wrong units and would have had to be converted prior to using it to calculate the speed of sound.
@@ VI.1 The role of prior knowledge and trustworthiness
The importance of prior knowledge about the specific area to successfully identify the validity of answers given by ChatGPT is further strengthened by examining an interaction during the lab by Group 4. Although they were one of the groups with second-year students taking Physics 1 and thus having access to very little prior knowledge of the physics of sound, they were identified as performing well during the lab since they were able to identify what data they should collect and perform the analysis without issues. However, they ended up struggling towards the end of the lab session when they were tasked to discuss how acoustic levitation works. Due to their limited to non-existent knowledge of wave mechanics and the physics of sound, they had few tools available to them to challenge what answers ChatGPT gave them. As a result, despite Group 4 demonstrating an understanding of the need to pose not just one question but many to try to identify if the answers provided by ChatGPT made sense or not, they were ultimately led down a path toward a wrong explanation that it is mainly the lower speaker that supplies a force by pushing on the styrofoam particle and thus counteracting gravity.
Finally, in most other interactions with ChatGPT during the lab session, the groups seemed to take the answers provided by ChatGPT as simply being the truth. This may be connected to how the students, at times, anthropomorphized ChatGPT during the interviews. Group 3 stated that they might not fully trust the answers if ChatGPT did not provide any sources or citations, and Group 5 stated that ChatGPT can be trusted since it has knowledge about previous research results. In general, during the group interviews before and after the lab, when asked about the capabilities of ChatGPT, students were quick to use words such as knowledge and talk less about it as some tool and more as a person with much knowledge. During the interviews, it was mainly groups 6 and 7 that stood by the argumentation that what ChatGPT can answer and not reliably is dictated by what data was part of its training, which indicates that they are more cognizant that GenAI tools do not have knowledge and reasoning skills similar to humans.
VI.2 Utilizing ChatGPT as a lab partner or a source of information
@@
Concerning RQ2 and RQ3, they focus more on how students interacted with ChatGPT during the lab to overcome specific hurdles. Previous studies provide little information about its capabilities to function as a lab partner. What it does show, however, is that depending on the type of questions and problems students face, it might be able to provide reliable answers, as it can be good at solving subject-specific problems [8, 9]. Further, it has also been identified that GenAI tools, like ChatGPT, can reproduce language and provide answers relevant to tackling problems in laboratory settings [54]. Our findings indicate similar conclusions. In RQ2, How do high school students actively engage with ChatGPT to solve problems during a lab investigating the speed of sound using acoustic levitation, Fig. 3 shows that students tend to make use of ChatGPT after having done all or most of the data collection. This indicates that despite having little to no knowledge about the phenomenon under study, the groups had little problem collecting data using the experimental setup.
@@
During the group interviews, students stressed that a major use case of ChatGPT is to get answers to conceptual questions or to clarify problems or instructions, and that it made little or no sense in asking about lab equipment. This sentiment is in agreement with how the groups did make use of ChatGPT during the lab session: during the analysis and to get help with the discussion questions about acoustic levitation. Here, the primary beneficial use case for the groups was related to getting help with identifying what equations to make use of, as well as getting an introduction to acoustic levitation. A noteworthy special case was Group 5, who “outsourced” the analysis to ChatGPT after failing to understand what to do with the data by simply asking what the sound speed was at the room’s current temperature. During the analysis, it is important to stress that the lab instructions were intentionally vague in that no equations were provided. Instead, the assumption was that students would utilize a combination of ChatGPT and their formula book to identify how to make use of the data to get the speed of sound using the equation $v=f\lambda$ and compare that to the speed of sound they get when using an equation that included temperature dependence. Additionally, if the groups had time, they were tasked to discuss questions related to acoustic levitation, including the underlying principles of acoustic levitation. As such, it was expected that all groups, regardless of their natural intention to use or not use ChatGPT during the study, were incentivized to at least try to use it during the analysis and discussion part.
VI.3 GenAI as a tool to develop problem-solving skills in laboratory settings
@@
Overall, the findings presented in the current study demonstrate fruitful and problematic interactions with ChatGPT in terms of aiding students in solving problems or helping them understand the phenomenon of acoustic levitation. Groups 4 and 6 were both able to use ChatGPT to aid them during the analysis in that Group 4 successfully utilized it to aid them in reasoning about whether the wavelength they got from the data collection was reasonable. Further, both groups showed proficiency in discussing how to identify if the equations given to them from ChatGPT were reliable by drawing upon their prior knowledge of dimensional analysis. Contrary to these, groups 1, 2, 3, and 5 all demonstrated that lacking knowledge about the relevant physics and how GenAI tools like ChatGPT “reasons” or has “knowledge”, having access to such tools in a classroom setting, including laboratories, might end up disadvantageous for the students learning process.
In this contrast between groups 4 and 6, and groups 1, 2, 3, and 5, we can see a situation where ChatGPT can serve as a tool to help reduce the importance of prior content knowledge during lab sessions, which teachers can utilize to help their students develop certain capabilities. Namely, that groups 4 and 6, despite significant differences regarding relevant knowledge of the underlying physics, both were able to make use of ChatGPT to assist them in moving forward with the data analysis. What instead separated these groups from 1, 2, 3, and 5, was their assumptions and knowledge about how GenAI tools function and “reasons.” The kind of interactions groups 4 and 6 demonstrated during the analysis provided them with new tools on how to approach problem-solving in a laboratory context. Drawing upon variation theory, we can identify an alternative approach to the lab session that could allow the struggling groups to succeed more similarly to groups 4 and 6. If an object of learning is chosen related to utilizing GenAI tools in problem-solving during laboratory work, a potential critical aspect may be related to discerning how current GenAI tools’ reasoning differs from that of humans. Here, the argument for the critical aspect is connected to the identified stark difference between groups 4, 6, and 7 and the other groups when it came to assumptions about how GenAI tools work. Such a focus would also open up for more active scaffolding regarding steering the students’ focus towards critical aspects [36]. This would include rather heavily aiding students with content-specific conceptual understanding, leaving them to instead focus on problem-solving and GenAI-related aspects. Finally, this shift in focus towards more general skills has been identified in previous work in PER using variation theory, as exemplified by [42] on how students develop skills related to developing internal idealized models and use them to develop conceptual understanding in the physics laboratory (referred to as being able to modelize). It is also a focus that has widespread acceptance in the current PER literature as a means towards getting a more expert-like view of laboratory work in physics (e.g., [55, 56, 57, 58]).
Continuing, we can connect the suggested change in focus regarding the object of learning to RQ3, How can interacting with ChatGPT during the lab aid students in solving problems and with conceptual understanding. With the aim of having students develop new skills using GenAI when solving problems during laboratory work, a crucial part of their scaffolding during the lab would revolve around probing the students on their conceptual understanding of the phenomenon studied during the lab by actively interacting with the groups during the lab. We base this suggestion on that our findings show instances of ChatGPT providing information to the students about the underlying physics, which is wrong or misleading. An example of such an instance is seen in the interaction Carl in Group 3 has with ChatGPT. There, the analogy of keeping a ball in the air by blowing on it from below reinforced the notion that the main contributor to the phenomenon seen during the lab was due to the bottom speaker pushing air particles on the styrofoam particle. If a teacher were present and identified Carl’s acceptance of this analogy, they could intervene and provide an explanation to correct the problems with the analogy. Thus, Carl and the rest of Group 3 would be back towards the intended object of learning instead of having their awareness shifted towards aspects related to a more concept-focused object of learning.
VI.4 The ongoing problem with GenAI and hallucinations
Becoming better at addressing the issue of getting problematic and made-up answers from GenAI tools in the classroom relates to the growing need for AI literacy in education, for students and teachers alike [21, 26, 59]. As is evident from our findings, and in agreement with previous work [60], students with a combination of less knowledge of the subject (here, physics) and not being familiar with GenAI tools or have little to no knowledge about how they function struggle with making use of these new tools. Proposed solutions from the literature involve training students and teachers in prompt engineering [21, 59, 61]. This proposal is further supported by our findings, where groups 4 and 6 utilized their understanding and assumptions about how ChatGPT functions to help them distinguish between reliable and unreliable information in the answers provided by the chatbot. From their interviews and observations of their lab sessions, it became evident that their assumptions led them to ascribe a less human-like awareness in how ChatGPT generated responses.
Developing skills related to prompt engineering can be a fruitful way to give students and teachers a greater understanding of not only how to extract more useful information from GenAI tools, but also to better comprehend changes in limitations over time between and within models [21, 59]. For educators, an important step in learning how the behavior of GenAI tools varies over time is for researchers to continuously monitor how various models perform, with and without proper prompt engineering and further fine-tuning. To that end, the current study has explored how largely untrained students can and do make use of a free version of ChatGPT, GPT-3.5 at the time of data collection, when solving problems during a physics lab. The findings and discussions presented here can benefit educators and researchers in gaining a deeper understanding of the naïve and partially spontaneous interaction between students and GenAI, particularly in identifying fruitful ways of interaction and problematic instances of overreliance on ChatGPT as a substitute for or absolute source of information.
The use of GPT-3.5, rather than, for example, GPT-4o, has profound implications in terms of the type of misinformation presented by ChatGPT, as well as how it responds to questions in this study. Additionally, it is reasonable to assume that the way ChatGPT behaves has an effect on how participants in this study rely on the answers and find ChatGPT helpful. However, subsequent releases of GenAI tools still struggle with the problem of hallucinations, albeit less frequently, and it is a problem that may be inherent to the way large language models work [62]. This implies that although the specific problems ChatGPT exhibits in this study may not be present if replicating the study today, we would expect similar conclusions, but perhaps that more groups may have been able to successfully use ChatGPT, related to when GenAI tools can be useful and why when considering what is to be learned.
VII Conclusion and outlook
@@
In the current study, we address the research gap related to the need of exploring how GenAI tools, specifically the free version ChatGPT at the time of the data collection, could be used in the educational physics laboratory. Our findings are in agreement with previous research regarding the importance of educating students about the capabilities and limitations of AI tools, as well as gaining some level of understanding about how they are developed and “think.” Additionally, in our analysis, we have explored a novel way of conceptualizing the enacted object of learning from a student perspective, and its connection to the structure of awareness. This allowed us to gain further insight into how GenAI tools can be utilized in laboratory settings to develop students’ problem-solving skills.
In STEM education specifically, there is a need to further explore and develop educational strategies and material that can help students and teachers better understand how to best utilize and think about the role of GenAI in education, not the least in laboratory settings. As these tools continue to be developed, their capabilities to simulate a trustworthy chain of reasoning will improve, furthering the need to carefully consider what implications interactions with such tools may have when it comes to learning concepts in disciplinary-accepted ways. Here, the role and knowledge of the teacher about AI is of utmost importance. Finally, we have identified that there is a potential to make use of GenAI tools like ChatGPT in the physics laboratory as a lab partner, but that the usefulness looks to be dependent on students knowledge about the phenomenon under study as well as their knowledge about how GenAI tools function. Furthermore, the specific ways in which such a tool is best employed in the physics laboratory require more research and will also be dictated by the future development of such tools.
@@ Acknowledgements.
Sebastian: Conceptualization, Data curation (lead), Formal analysis, Investigation (lead), Methodology, Project administration, Resources (lab design; supporting), Validation (lead), Visualization (lead), Writing - original draft, Writing - review & editing (lead).
Andreas: Investigation (supporting), Resources (lab design; lead), Validation (supporting), Visualization (supporting), Writing - original draft (supporting), Writing - review & editing (supporting).
Jonas: Data curation (supporting), Investigation (supporting), Supervision, Validation (supporting), Writing - review & editing (supporting).
Appendix A Lab Manual
Speed of sound and wavelength
Aim
To explore acoustic levitation using LeviLab and determine the speed of sound in air.
Equipment
  • LeviLab (two ultrasonic speakers mounted on a caliper, a function generator, an amplifier, and a thermometer)
  • Lab stand equipment
  • Voltage supply (DC, 20 V)
  • Polystyrene particles
  • ChatGPT
Experiment
  1. 1.
    Information, assembly, and operation of LeviLab
    • Place LeviLab on a flat surface.
    • Stabilize LeviLab using the lab stand equipment.
    • Vary the loudspeaker distance using the caliper.
    • The emitted sound frequency is 40 kHz.
  2. 2.
    Calculation of the speed of sound in air
    • Verify that the power supply is on.
    • Place a polystyrene particle on the lower speaker.
    • Gradually lower the upper loudspeaker and identify points where the particle appears to ‘jump up’ and start levitating. Does increasing the distance between the speakers produce the same effect?
    • Place the loudspeakers at a distance from each other so that as the loudspeaker distance approaches zero, the particle has jumped 7-10 times. Here is a closer description of a measurement:
      • Carefully lower the top loudspeaker. Stop exactly at the position where the polystyrene particle jumps up and starts to levitate.
      • Use the caliper to record the distance between the loudspeakers.
      • Record the temperature of the ambient air using the thermometer.
  3. 3.
    Analysis
    • Calculate the wavelength
      • Label your measurements such that 1 represents the shortest speaker spacing. This is called an ordinal number.
      • Conduct a linear regression analysis with loudspeaker spacing as the dependent variable and the ordinal number as the independent variable. The slope of the regression line corresponds to the wavelength.
    • Calculate the speed of sound using a formula involving wavelength and frequency.
    • Compare your result with the speed of sound as determined by the temperature reading. This is achieved by applying a linear model that expresses the relationship between the speed of sound and the temperature in degrees Celsius.
Discuss in the group
  1. 1.
    What are the fundamental principles of acoustic levitation that make it possible for sound waves to balance the gravitational force on an object? Construct a descriptive figure.
  2. 2.
    Discuss possible sources of error and how to minimize them to calculate more reliable speed of sound values.
  3. 3.
    Explain why it is important to measure the ambient temperature when calculating the speed of sound.
  4. 4.
    Discuss potential applications of acoustic levitation in situations outside school.
  5. 5.
    Explore and describe other applications where standing waves are used.